can you solve it without calculus?
Cl = Cз = 2 и 4г, 02 = C = 6 цE у?? / – 12 / се С2
Answer #1

by 49 ใ น 9 ใ3 น ? +-+-+ = 0 + C2 V 442 - 12x 8 x 2 UC [2+8 : 3,248 5 สี Ky2 40 q = 57obul 38, *-* - 59-646similarly. v. 1 - 2 20 p = Ozly v t estly 12808 -X12 8+12 Q = 57.646 q=0 = 57.646 Soi q =%2=42=24 = 57.646 92= 6 +62 = zouf [А Allolto С. Сач q= 920C BY XV 12 tozu - 20x20 x 12 20120 9 = 120 ut VAO = t = 120 16 = 6V ZVOR 192 zou So 19 = 94 = 12x6 =42